OpenStudy (anonymous):

At a town fair, 9 tickets cost $12. How many tickets could you expect to receive for $8, assuming there is no discount for buying more tickets? a] 4 b] 5 c] 6 d] 7

OpenStudy (anonymous):

6 tickets

OpenStudy (anonymous):

thats wrong

OpenStudy (anonymous):

9/12 = x /8

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

9*8 = 12x divide by 12

OpenStudy (anonymous):

x = 6 tickets

OpenStudy (anonymous):

12/9 will give you a number that represents dollars per ticket. $1.33. So 1 ticket costs $1.33. So $8=$1.33x Solve for x.